### v1

@@ Abstract
The ocean is teeming with a myriad of submillimeter-sized invertebrate planktonic larvae [1], which thrive in a viscous fluid environment [2]. Many of them rely on ciliary beating to generate fluid flows for locomotion and feeding [2, 3]. Their forms, local morphologies, and ciliation patterns exhibit remarkable diversity [4], producing intricate and dynamic three-dimensional (3D) flows that are notoriously difficult to characterize in laboratory settings. Traditional microscopic imaging techniques typically involve gently squeeze-confining the soft larvae between a glass slide and cover slip, to study their flows in quasi-two-dimensions (2D) [5]. However, a comprehensive hydrodynamic framework for the low-to-intermediate Reynolds number (1) flows at the larval scale in quasi-2D confinement—particularly in light of their complex forms—has remained elusive. Here, we demonstrate that vortices around larvae proliferate with increasing confinement, and elucidate the underlying physical mechanism. We experimentally quantify confinement-induced flows in larvae of sea stars and sea urchins. The flows exhibited strikingly universal patterns: under weak confinement, all larvae generated two vortices, whereas under strong confinement, multiple vortices appeared. The experimental observations were well captured by a low Reynolds number theoretical model based on superposition of confined Stokeslets [6, 7]. Building on experiments and theory, we developed a comprehensive framework for confinement-induced larval flows, which suggests that the vorticity dynamics are primarily determined by local morphological features, rather than the larval forms. Our work provides fundamental insights into the form-functional relationships between larval morphology and flow generation. Our findings are broadly applicable to understanding fluid flows generated by a wide range of ciliated organisms with complex forms and morphologies, from micro- to milli-length-scales.
@@
Biomechanics, Fluid dynamics, Marine invertebrate larvae, Confinement, Ciliary flowsDiverse and exotic life forms inhabit the vast oceans and hold many secrets to life that are waiting to be unraveled [1]. The biomechanics of marine animals and their interactions with the fluid environment are fundamental drivers of ecological and evolutionary processes in the ocean. Marine species generate fluid flows for a variety of purposes, including locomotion, feeding, predator evasion, reproduction, and development [8, 2, 9, 10, 11]. Many benthic marine invertebrates, such as echinoderms (e.g. sea stars and sea urchins) undergo an indirect mode of development that involves a pelagic larval stage with a form completely different from their adult form. During this larval stage, these organisms swim and feed in the water column for weeks or even months, before metamorphosing into their adult shape. The majority of these larvae are soft-bodied planktonic free-swimmers that utilize ciliary beating to create fluid flows that are used for both swimming and feeding [3, 5, 12]. In these larvae, large numbers of motile cilia are arranged in ‘ciliary bands’ that are intricately embedded in their body morphologies, and enable them to generate unexpectedly rich fluid-flow patterns. So far, sea star larvae have been shown to generate a fascinating array of multiple counter-rotating vortices around them [5, 13]. Further, the diversity of ciliated larval forms is staggering, with remarkable variations in body morphology and ciliation patterns [4]. Therefore, an open question is whether the differences in forms and local morphologies between ciliated larvae give rise to distinct flow patterns? We will investigate this question for the first time by examining morphologically distinct larvae from two closely related echinoderms: the sea star (Patiria miniata) and sea urchin (Lytechinus variegatus). Specifically, we will investigate two developmental stages (Early-stage and Late-stage) of sea star larvae, and one stage of sea urchin larvae. The sea star larvae have an oval/ellipsoid form, with bilateral and anterior-posterior symmetry; the Early-stage bipinnaria larvae have a simpler morphology compared to the more complex-shaped Late-stage brachioleria larvae, which feature short arms. In stark contrast, the echinopluteus sea urchin larvae have a cone-shaped form, with rigid, extended skeletal arms and exhibit bilateral symmetry but lack anterior-posterior symmetry. In this study, we aim to experimentally quantify and model the flow patterns generated by these three distinct larval types, seeking to uncover the fundamental relationships between the different forms of marine invertebrate larvae and their hydrodynamic signatures.
@@
The tiny submillimeter-size larvae reside in a viscous fluid environment in the ocean, where viscous forces dominate over inertial forces. This balance is quantified by the Reynolds number (), which represents the ratio of inertial to viscous forces, and is defined as , where , , and are the characteristic velocity, length scale, and kinematic viscosity of water, respectively. For sea star and sea urchin larvae swimming in seawater, 0.1–0.9 (Table 1, Supplementary Information), placing them within a low-to-intermediate Reynolds number regime, where although viscous effects are prominent, inertial effects are also starting to play a role. Low-to-intermediate Reynolds number flows represent a relatively under-explored regime in fluid dynamics, as the majority of the existing literature on biological fluid mechanics focuses on either low Reynolds number flows [2, 3, 14] or intermediate to high Reynolds number flows [8, 9]. In this study, we present new experimental results at 0.1–0.9, and demonstrate that theoretical models based on low Reynolds number assumptions are able to provide a qualitatively accurate description of the observed phenomena.
@@
The majority of aquatic microorganisms, ranging from micro- to milli-scale, swim at low-to-intermediate Reynolds numbers freely in three dimensions (3D) within open water bodies such as lakes and oceans. However, these microorganisms frequently encounter rigid boundaries or surfaces due to interactions with other living organisms or non-living structures, such as the air-water interface or other solid matter, including substrates at the bottom. Microorganism-boundary interactions have significant biological implications in various natural contexts. For example, sessile, suspension-feeding ciliated organisms like Vorticella [15, 16], Stentor [17], and Coral [18, 19] need to optimize nutrient availability in their immediate surroundings. Similarly, many organisms rely on internal ciliary-driven flows within ducts to regulate vital physiological processes [20]. The marine invertebrate larvae examined in this study likely come across rigid boundaries primarily during terminal encounters or in laboratory settings. Nevertheless, they serve as an excellent model system for investigating the hydrodynamic effects of microorganism-boundary interactions. Confinement-induced fluid dynamical phenomena are often found to be highly non-intuitive and non-trivial, garnering considerable interest in the field of biological fluid dynamics, particularly at low Reynolds numbers [6, 21, 22, 15, 16, 23, 5, 24, 25, 26, 7, 27, 28, 17, 29, 30]. In the present work, we investigate confinement effects in low-to-intermediate Reynolds number flows for the first time.
@@
The complex 3D flow-fields generated by ciliated marine invertebrate larvae pose significant challenges for quantification in laboratory-based experimental studies, even though there have been recent advancements [31]. The straightforward traditional microscopic imaging methods, which confine larvae between a glass slide and a cover slip, impose a quasi-two-dimensional (2D) confinement. So far, this conventional quasi-2D approach has been instrumental in revealing phenomena such as the multiple vortex arrays generated by sea star bipinnaria larvae [5]. However, alternative confinement techniques, such as tethering larvae [24], fail to produce the same multiple vortex patterns [5, 25, 13]. This discrepancy raises fundamental questions about the impact of quasi-2D confinement on the flow fields of 3D micro-swimmers. Here, we resolve this conundrum and address several longstanding questions that have puzzled biological fluid mechanicians for decades: How does quasi-2D confinement alter the fluid flow fields of a 3D micro-swimmer? Do variations in geometric form and local morphology of ciliated larval micro-swimmers give rise to different flow patterns under quasi-2D confinement? Is there a unified theoretical hydrodynamic framework that can capture these quasi-2D confinement effects at low-to-intermediate Reynolds numbers?
@@
We address these open questions in physics and biology by investigating the effects of squeeze-confinement on the low-to-intermediate Reynolds number flows generated by three distinct forms of ciliated marine invertebrate larvae of echinoderms: (i) Early-stage sea star, (ii) Late-stage sea star, and (iii) Sea urchin (Fig. 1). We systematically vary the confinement parameter—the height between the glass slide and cover slip ()—and quantify changes in the vorticity dynamics. We quantitatively compare our experimental results with a theoretical low Reynolds number-based model, and develop a universal hydrodynamic framework for confinement flows.
@@
Using live darkfield time-lapse microscopy experiments, we observed that the flow-fields generated by the three distinct larvae varied depending on the degree of squeeze-confinement (Methods, Fig. 1). The larvae were confined by gently trapping or squeezing them in between a glass slide and a cover slip that were separated by a precisely known height ’H’, i.e. the squeeze-confinement parameter. This confinement parameter was varied over a broad range for the three larvae – from ’weak confinement’ (low squeeze/large H) to ’strong confinement’ (high squeeze/small H) and their resultant flow-fields were quantified using standard techniques from experimental fluid dynamics (Fig. 1, Methods). The flow visualizations in Fig. 1 (a-f) were carried out by post-processing of the time-lapse images using Flowtrace in ImageJ [32] (left panels), and Particle Image Velocimetry (PIV) in MATLAB [33] (right panels) (Methods). We observed that weak confinement (low squeeze/large H) resulted in two vortices in all the three types of larvae (Fig. 1 a,c,e) (Video 1). When the confinement was increased by bringing the glass slide and cover slip closer, first reaching ’moderate’ confinement conditions (medium squeeze/medium H) (Extended Data Fig. 1) and then going to ’strong’ confinement conditions (high squeeze/small H), all the larvae generated more than two vortices (Fig. 1 b,d,f) (Video 2). In our quasi-2D experiments, the thin imaging z-plane was focused close to the mid-plane () located in between the bottom (glass slide) and top (cover slip) boundaries. We found that there was minimal flow field variation in the z-planes above and below the mid-plane (Extended Data Fig. 2). Hence our experiments captured the dominant and representative quasi-2D flow fields generated by the larvae.
@@
To account for the size variation among the three types of larvae, we divide the confinement height ’H’ length scale with the larval depth dimension ’c’, to arrive at a normalized squeeze-confinement parameter (Supplementary Information). This non-dimensional confinement parameter provides a standardized quantification of the degree of squeeze-confinement, and hence enables comparison of flows between the three larval types. We found that weak confinement (squeeze parameter, ) always resulted in a fixed number of vortices (two) being generated by the larvae, irrespective of their form and local morphology (Fig. 1 a,c,e). However, when the strength of confinement was increased (squeeze parameter, ), the total number of vortices generated varied depending on the degree of confinement and the larval form and local morphology (Fig. 1 b,d,f). Under strong confinement () conditions, the simple-shaped Early-stage sea star larva developed four total vortices, i.e. two symmetric pairs (Fig. 1 b). Although the complex-shaped Late-stage sea star and Sea urchin larvae developed a total of six vortices under strong confinement () (Fig. 1 d,f), the vortex locations were not similar since these two larvae have completely different forms and complex morphologies. Meanwhile, in the moderate confinement regime (), the three different types of larvae generated more than two vortices (Extended Data Fig. 1), but less than the maximum number they generated under strong confinement (Fig. 1 b,d,f). Taken together, these results suggest that both the confinement parameter and larval morphology play an important role in determining vortex dynamics under confinement.
@@
Further quantification of the confinement-induced fluid flows reveals several interesting common trends. The total number of vortices generated over a range of confinement parameters (0.1–1) was determined using a large number of samples from our experiments (Fig. 1 h) (Extended Data Fig. 3 a,b) (Table 1, Supplementary Information). Irrespective of the larval form and morphology, we observed that weak confinement () always resulted in 2 vortices. However, when the degree of squeeze-confinement was increased (), the number of vortices increased with an almost linear trend (inferred via linear fits) with slightly different slopes for all the three larvae (Fig. 1 h). Under the strongest squeeze-confinement conditions () (Fig. 1 h), both the morphologically complex Sea urchin larvae with extended arms, and the morphologically complex Late-stage sea star larvae, developed a maximum of nine vortices (Fig. S1, Supplementary Information). However, the morphologically less complex Early-stage sea star larvae only generated a maximum of six vortices (Fig. S1, Supplementary Information). Overall, our results reveal that all the three types of larvae generate two vortices under weak confinement, and the number of vortices increase with confinement following an almost linear trend despite significant variations in their form and local body morphology (Fig. 1 h). From a physical perspective, these results suggest that more number of vortices are generated due to increased frictional effects when the top (cover slip) and bottom boundaries (glass slide) come closer to each other. Hence, we believe that this general trend of number of vortices increasing with increased confinement is universal and can be applicable to a wide majority of marine ciliated larvae at comparable sizes. Also, the intricate details of the larva’s local body morphology seems to determine the slope of the linear trend, with the most complex larval morphologies generating the largest number of vortices under strong confinement.
@@
In addition to the total number of vortices generated, the larval flow-fields under different confinement conditions also showed systematic variations in other flow metrics such as the mean flow speeds generated (Fig. 1 i), and center locations of vortices (Fig. 1 j) (Methods). We observed that the mean flow speeds were maximum under weak confinement (), and the speeds decreased almost linearly with increasing confinement () (Fig. 1 i, Extended Data Fig. 3 c). This trend can be attributed to an increase of frictional effects from the walls due to increased confinement (), leading to suppression of the strength of fluid flows (speeds) generated by the larvae. We also observed that Sea urchin larvae generate fluid flows of higher mean speeds compared to the sea star larvae under weak confinement (Fig. 1 i), and this is also reflected in the different slopes of linear fits for the three larvae. Next, our quantification of the distances from vortex centers to the larval body surface, also showed a maximum value under weak confinement (), and a similar decreasing linear trend with increased confinement () (Fig. 1 j, Extended Data Fig. 3 d). This result implies that the vortices come closer to the larval body surface with increase in confinement and vice-versa. Finally, we also quantified the mean vorticity and circulation generated by the three larvae, and once again found a similar trend of linear decrease in values with increased confinement (Extended Data Fig. 2 e,f). Hence, all our experimental flow quantification results support our physical interpretation of increasing frictional effects from the walls due to increased confinement. Overall, these experimental results suggest the existence of an underlying hydrodynamic mechanism that determines the vortical flow patterns under different confinement conditions.
@@
In order to investigate the hydrodynamics of the confinement-induced flow fields in the larvae, we quantified the vorticity in the experimental data and developed a theoretical model that closely captures the experimental flow-fields. The hydrodynamic regime is quantified by the Reynolds number (), which is a ratio of inertial to viscous forces, given by , where and are the characteristic velocity and length scales, and is the kinematic viscosity of the fluid. The for the three types of marine larvae studied here are in the range (Methods, Supplementary Information, Table 1), which corresponds to a low-to-intermediate range where the flows are more dominated by viscous forces compared to inertial forces. Hence, we adopted a recently developed low hydrodynamics model [7] to quantitatively compare the experimental results. The model based on the quasi-2D Stokes equation for low flows consists of superposition of pairs of Stokeslets between two parallel flat plates, and explicitly takes into account the swimmer (larva) force F, and the chamber height H (squeeze-confinement parameter) [7, 6]. The final velocity field expressed in polar coordinates () is given by:
@@
The comparison of the flow-field results obtained from PIV-analysis of our experimental data [33, 34] and our theoretical model are shown in Fig. 2. In weak confinement experiments (), all the larvae – Early-stage and Late-stage sea star, Sea urchin – in spite of their distinct forms and local morphological variations, generated one dominant (large) pair of counter-rotating vortices (Pair 1) (Fig. 2 a, c, e, left panels).
@@
In the oval-shaped Early-stage sea star larva and more complex-shaped Late-stage sea star larva under weak confinement, the pair of vortices are roughly circular-shaped, and their centers are located towards the back-end of the body (Fig. 2 a, c, left panels). The theoretical model utilized two pairs of Stokeslets for Early-stage sea star and three pairs of Stokeslets for the Late-stage sea star larvae under the same experimental confinement conditions (), and physical location around the body surface (Methods, Supplementary Information), and showed excellent agreement with the experiments for both types of sea star larvae (Fig. 2 a, c, right panels) (Extended Data Fig. 4 a,d). It must be noted that in the theoretical model, each Stokeslet generates 1 pair of vortices. However, the resultant velocity field due to the superposition of multiple Stokeslets gave rise to a flow field consisting of two vortices under weak confinement, closely matching the experiments (Extended Data Fig. 4, Methods, Supplementary Information).
@@
However, in Sea urchin larvae, weak confinement resulted in vortices that are highly elongated in an ellipsoidal shape and distributed over the sides of the arms. In addition, the vortex centers are located near the tip of the extended arms (Fig. 2 e left panel). The theoretical model here utilized two pairs of Stokeslets under the same experimental confinement conditions () and physical location at the tip of the extended arms (Methods, Supplementary Information), and showed a flow-field comparable in magnitude with the experiments, but not in the number of vortices generated (Fig. 2f right panel). Compared to the sea stars, the Sea urchin larval morphology is much more complex, with with extended rigid arms. In this developmental stage, the Sea urchin larva also had two small rudimentary arms near its oral surface, and these could also potentially influence the observed flow- fields. For simplicity, the theoretical model does not account for the presence of these small arms. Hence, the theoretical flow-fields are slightly different from that of experiments. Overall, each of the experimental larval vortices under weak confinement can be modeled theoretically with the superposition of multiple Stokeslets, and the experimental two-vortex larval flow-field could be captured accurately.
@@
Under strong confinement conditions (), the local morphological variations between the three larvae play an important role in the number of vortices generated (Fig. 2 b, d, f, left panels). While the Early-stage sea star generates two pairs of vortices (Pairs 1,2) (Fig. 2 b), both the Late-stage sea star and Sea urchin larvae generate three pairs of vortices (Pairs 1,2,3) (Fig. 2 d, f). Hence, strong confinement promotes the development of additional pairs of vortices in all the larvae (Fig. 2 b, d, f). The Early-stage sea star larva has a oval/ellipsoid-shaped form, and although it does not yet have arms protruding outside the body, there is a sharp convex curvature change in the region where arms will be formed in the future. The theoretical model with two pairs of Stokeslets placed at exactly the region corresponding to the convex curvature change shows an excellent agreement with the experiments (Fig. 2 b, right panel).

### v2

@@ Abstract
The ocean is teeming with a myriad of submillimeter-sized invertebrate planktonic larvae [1], which thrive in a viscous fluid environment [2]. Many of them rely on ciliary beating to generate fluid flows for locomotion and feeding [2, 3]. Their larval forms, local morphologies, and ciliation patterns exhibit remarkable diversity [4], producing intricate and dynamic three-dimensional (3D) flows that are notoriously difficult to characterize in laboratory settings. Traditional microscopic imaging techniques typically involve gently squeeze-confining the soft larvae between a glass slide and cover slip to study their flows in quasi-two-dimensions (2D) [5]. However, a comprehensive hydrodynamic framework for the low-to-intermediate Reynolds number (1) flows at the larval scale in quasi-2D confinement — particularly in light of their complex forms — has remained elusive. Here, we demonstrate that vortices around larvae proliferate with increasing confinement and illuminate the underlying physical mechanism. We experimentally quantify confinement-induced flows in larvae of sea stars and sea urchins. The flows exhibited strikingly universal patterns: under weak confinement, all larvae generated two vortices, whereas under strong confinement, the number of generated vortices significantly increased. The experimental observations were well captured by a low Reynolds number theoretical model based on the superposition of confined Stokeslets [6, 7]. Building on experiments and theory, we developed a comprehensive framework for confinement-induced larval flows, which suggests that vorticity dynamics are primarily determined by local morphological features, rather than solely the body plan. Our work provides fundamental insights into the form-functional relationships between larval morphology and flow generation. Our findings are broadly applicable to understanding fluid flows generated by a wide range of ciliated organisms with complex forms and morphologies, from micro- to milli-length-scales.
@@
Biomechanics, Fluid dynamics, Marine invertebrate larvae, Confinement, Ciliary flowsDiverse and exotic life forms inhabit the vast oceans and hold many secrets to life that are waiting to be unraveled [1]. The biomechanics of marine organisms and their interactions with the fluid environment are fundamental drivers of ecological and evolutionary processes in the ocean. Marine species generate fluid flows for a variety of purposes, including locomotion, feeding, predator evasion, reproduction, and development [8, 2, 9, 10, 11]. Many benthic marine invertebrates, such as echinoderms (e.g. sea stars and sea urchins) undergo an indirect mode of development that involves a pelagic larval stage with a body plan completely different from their adult form. During this larval stage, these organisms swim and feed in the water column for weeks or even months before metamorphosing into their adult shape. The majority of these larvae are soft-bodied, planktonic, free-swimmers that utilize ciliary beating to create fluid flows for both swimming and feeding [3, 5, 12]. In these larvae, large numbers of motile cilia are arranged in ciliary bands embedded in their body morphologies enabling them to generate unexpectedly rich fluid-flow patterns. So far, sea star larvae have been shown to generate a fascinating array of multiple counter-rotating vortices [5, 13]. Further, the diversity of ciliated larval forms is staggering, with remarkable variations in body morphology and ciliation patterns [4]. Therefore, an open question is whether the differences in organism body plans and/or local morphologies between ciliated larvae give rise to distinct flow patterns. We will investigate this question for the first time by examining morphologically distinct larvae from two closely related echinoderms: the sea star (Patiria miniata) and sea urchin (Lytechinus variegatus). Specifically, we will investigate two developmental stages (early stage and late stage) of sea star larvae, and one stage of sea urchin larvae. Sea star larvae have a smooth oval/ellipsoid body plan, with bilateral and anterior-posterior symmetry; the early stage bipinnaria larvae have a simpler body plan compared to the more complex-shaped late stage brachioleria larvae, which feature local morphological variations. In stark contrast, the echinopluteus sea urchin larvae have a cone-shaped body plan, with rigid, extended skeletal arms and exhibit bilateral symmetry but lack anterior-posterior symmetry. In this study, we aim to experimentally quantify and model the flow patterns generated by these three distinct larval types, to uncover the fundamental relationships between the different forms of marine invertebrate larvae and their hydrodynamic signatures.
@@
The tiny, sub-millimeter larvae reside in a viscous fluid environment in the ocean, where viscous forces dominate over inertial forces. This balance is quantified by the Reynolds number (), which represents the ratio of inertial to viscous forces, and is defined as , where , , and are the characteristic velocity, length scale, and kinematic viscosity of water, respectively. For sea star and sea urchin larvae swimming in seawater, 0.1–0.9 (Table S1, Supplementary Information), placing them within a low-to-intermediate Reynolds number regime, where although viscous effects are prominent, inertial effects begin to play a role. Low-to-intermediate Reynolds number flows represent a relatively under-explored regime in fluid dynamics, as the majority of the existing literature on biological fluid mechanics focuses on either low Reynolds number flows ( to 0.1) [2, 14] or intermediate to high Reynolds number flows () [8, 9]. In this study, we present new experimental results at , and demonstrate that theoretical models based on low Reynolds number assumptions provide a qualitatively accurate description of the observed phenomena.
@@
The majority of aquatic microorganisms at low-to-intermediate Reynolds numbers swim freely in three dimensions (3D) within open water bodies such as lakes and oceans. However, these microorganisms frequently encounter rigid boundaries or surfaces due to interactions with other living organisms or non-living structures, such as the air-water interface or other solid matter, including substrates at the bottom. Microorganism-boundary interactions have important biological implications in various natural contexts. For example, sessile, suspension-feeding ciliated organisms like Vorticella [15, 16], Stentor [17], and coral [18, 19] need to optimize nutrient availability in their immediate surroundings since they are unable to swim away to find nutrient sources. Similarly, many animals rely on internal ciliary-driven flows within organs such as human airways, brain ventricles and reproductive tracts to regulate vital physiological processes [20].
The marine invertebrate larvae examined in this study likely come across rigid boundaries during terminal encounters or in laboratory settings. Nevertheless, they serve as an excellent model system for investigating the hydrodynamic effects of microorganism-boundary interactions. Confinement-induced fluid dynamical phenomena are often found to be highly non-intuitive and non-trivial, garnering considerable interest in the field of biological fluid dynamics, particularly at low Reynolds numbers [6, 21, 22, 15, 16, 23, 5, 24, 25, 26, 7, 27, 28, 17, 29, 30]. In the present work, we investigate confinement effects in low-to-intermediate Reynolds number flows for the first time.
@@
The complex 3D flow fields generated by ciliated marine invertebrate larvae pose many challenges to quantification in laboratory-based experimental studies, despite recent advancements [31]. The straightforward traditional microscopic imaging methods, which confine larvae between a glass slide and a cover slip, impose a quasi-two-dimensional (2D) confinement. So far, this conventional quasi-2D approach has been instrumental in revealing phenomena such as the multiple vortex arrays generated by sea star bipinnaria larvae [5]. However, alternative confinement techniques, such as tethering larvae [24], fail to produce the same multiple vortex patterns [5, 25, 13]. This discrepancy raises fundamental questions about the impact of quasi-2D confinement on the flow fields of 3D micro-swimmers. Here, we resolve this conundrum and address several longstanding questions that have puzzled biological fluid mechanicians for decades: How does quasi-2D confinement alter the fluid flow fields of a 3D micro-swimmer? Do variations in larval body plan and local morphology of ciliated larval micro-swimmers give rise to different flow patterns under quasi-2D confinement? Is there a unified theoretical hydrodynamic framework that can capture these quasi-2D confinement effects at low-to-intermediate Reynolds numbers?
@@
We address these open questions in physics and biology by investigating the effects of squeeze-confinement on the low-to-intermediate Reynolds number flows generated by three distinct forms of ciliated marine invertebrate larvae of echinoderms: (i) early stage sea star, (ii) late stage sea star, and (iii) sea urchin (Fig. 1). We systematically vary the height between the glass slide and cover slip () and quantify changes in the vorticity dynamics. We quantitatively compare our experimental results with a theoretical low Reynolds number-based model, and develop a universal hydrodynamic framework for confinement flows.
@@
Using live darkfield time-lapse microscopy experiments, we observed that flow fields generated by the three distinct larvae varied depending on the degree of squeeze-confinement (Methods, Fig. 1). The larvae were confined by gently trapping or squeezing them between a glass slide and a cover slip that were separated by a precisely known height, . This height was varied over a broad range for the three larvae – from ’weak confinement’ (low squeeze/large ) to ’strong confinement’ (high squeeze/small ) and their resultant flow fields were quantified using standard techniques from experimental fluid dynamics (Fig. 1, Methods). The flow visualizations in Fig. 1 (a-f) were carried out by post-processing of the time-lapse images using Flowtrace in ImageJ [32] (left panels), and Particle Image Velocimetry (PIV) in MATLAB [33] (right panels) (Methods). We observed that weak confinement (low squeeze/large ) resulted in two vortices in all the three types of larvae (Fig. 1a, c, e) (Video S1). When the confinement was increased by bringing the glass slide and cover slip closer, first reaching ’moderate’ confinement conditions (medium squeeze/medium ) (Extended Data Fig. 1) and then going to ’strong’ confinement conditions (high squeeze/small ), all the larvae generated more than two vortices (Fig. 1b, d, f) (Video S2). In our quasi-2D experiments, the thin imaging z-plane was focused close to the mid-plane () located in between the lower (glass slide) and upper (cover slip) boundaries. We found that there was minimal flow field variation in the z-planes above and below the mid-plane (Extended Data Fig. 2). Hence our experiments captured the dominant and representative quasi-2D flow fields generated by the larvae.
@@
To account for the size variation among the three types of larvae, we divide the confinement height () length scale with the larval depth dimension (), to arrive at a normalized squeeze-confinement parameter (Supplementary Information). This non-dimensional confinement parameter provides a standardized quantification of the degree of squeeze-confinement, and hence enables comparison of flows between the three larval types. Confinement strength can then be described categorically, using the same ranges for each larvae to describe similarities and differences between larvae body plans under strong (), moderate (), and weak () confinement conditions. We found that as confinement weakened ( increased), the number of generated vortices for all larvae, regardless of their body plan and local morphology, converged to 2 vortices (Fig. 1a, c, e, Supplementary Tables S8-S9). However, when confinement strength increased ( decreased), the total number of vortices generated varied depending on the confinement strength and larva body plan (Fig. 1b, d, f, Supplementary Tables S9-12). Under strong confinement () conditions, the simple-shaped, early stage sea star larva developed four total vortices, i.e., two symmetric pairs (Fig. 1b). Although the complex-shaped, late stage sea star and sea urchin larvae developed a total of six vortices under strong confinement () (Fig. 1d, f), the vortex locations were not similar since these two larvae have completely different body plans and local morphologies (Supplementary Tables S9, S12). Meanwhile, under moderate confinement conditions (), the three different types of larvae generated more than two vortices (Extended Data Fig. 1), but less than the maximum number they generated under strong confinement (Fig. 1b, d, f, Supplementary Tables S9-S11). Taken together, these results suggest that both the confinement parameter and larval morphology play an important role in determining vortex dynamics under confinement.
Regardless of the larval body plan and local morphology, we observed that as increased the number of vortices converged to 2 for all larval body plans (Fig. 1h). However, when the degree of squeeze-confinement increased ( decreases), the number of vortices increased (Fig. 1h, Supplementary Tables S8-11). Under the strongest observed squeeze-confinement conditions () (Fig. 1h), both the morphologically complex sea urchin larvae with extended arms, and the morphologically complex late stage sea star larvae developed a maximum of nine vortices, more than the simpler early stage sea star larvae, which only generated a maximum of six vortices (Fig. S1, Supplementary Information). In fact, there is a statistically significant difference between the number of vortices by sea urchin and early stage sea star larvae under strong confinement (, Supplementary Table S12).
@@
Overall, our results reveal that all three types of larvae generate two vortices under the weakest confinement (), and the number of vortices increase with confinement ( decreases) following an almost linear trend despite variations in their body plan and local morphology (Fig. 1h, Supplementary Table S8). From a physical perspective, these results suggest that a greater number of vortices are generated due to increased frictional effects when the upper (cover slip) and lower (glass slide) boundaries come closer to each other. Hence, we believe that this trend of increasing vortex number with increased confinement is universal and is applicable to a wide majority of marine ciliated larvae at comparable sizes. The number of vortices generated are also affected by the larva’s local morphology, with the most complex larval morphologies (sea urchin as the most complex, and early stage sea star as the least complex) generating significantly more vortices than simpler morphologies under strong confinement (Supplementary Tables S9-S10).
@@
In addition to the total number of vortices generated, the larval flow fields under different confinement conditions showed systematic variations in other flow metrics such as the mean flow speeds generated (Fig. 1i) and the distance of the vortices from the organism (Fig. 1j) (Methods, Supplementary Tables S9-12). We observed that the mean flow speeds were greatest under weak confinement () and the speeds decreased almost linearly with increasing confinement ( decreases) (Fig. 1i, Extended Data Fig. 3c, Supplementary Tables S10-11). This trend can likely be attributed to an increase of frictional effects from the walls due to increased confinement, leading to suppression of the strength of fluid flows (speeds) generated by the larvae. We also observed that sea urchin larvae generate fluid flows of higher mean speeds compared to the sea star larvae under weak confinement (Fig. 1i), however, these results are not significantly different than the sea star larvae, likely due to uneven sample size (Supplementary Table S12). Next, our quantification of the distances from vortex centers to the larval body surface also showed a maximum value under weak confinement (), and a similar decreasing linear trend with increased confinement ( decreased) (Fig. 1j, Extended Data Fig. 3d, Supplementary Tables S8-S10). This result implies that the vortices approach the larval body surface as confinement increased ( decreased). Finally, we also quantified the mean vorticity and circulation generated by the three larvae, and once again found a similar trend of linear decrease in values with increased confinement (Extended Data Fig. 2e, f, Supplementary Tables S8-S11). Hence, all our experimental flow quantification results support our physical interpretation of increasing frictional effects from the walls due to increased confinement. Overall, these experimental results suggest the existence of an underlying hydrodynamic mechanism that determines the vortical flow patterns under different confinement conditions.
@@
To investigate the hydrodynamics of confinement-induced flow fields around marine ciliated larvae, we quantified the vorticity in the experimental data and developed a theoretical model that closely matches the experimental flow fields. The for the three types of marine larvae studied here are in the range (Methods, Supplementary Table S1), which corresponds to a low-to-intermediate range where the flows are more dominated by viscous forces compared to inertial forces. Hence, we adopted a recently developed low hydrodynamics model [7] to quantitatively compare the experimental results. The model, based on the quasi-2D Stokes equation for low flows consists of superposition of pairs of Stokeslets between two parallel flat plates, and explicitly takes into account the swimmer (larva) force , and the chamber height (squeeze-confinement parameter) [7, 6]. The final velocity field expressed in polar coordinates () is given by:
@@
The comparison of the flow field results obtained from PIV-analysis of our experimental data [33, 34] and our theoretical model are shown in Fig. 2. In weak confinement experiments (), all the larvae, despite their distinct body plans and local morphological variations, generated one dominant (large) pair of counter-rotating vortices (Pair 1) (Fig. 2a, c, e, left panels).
@@
In the oval-shaped, early stage sea star larva and more complex-shaped, late stage sea star larva under weak confinement, the pair of vortices are roughly circular-shaped, and their centers are located towards the back-end of the body (Fig. 2a, c, left panels). The theoretical model utilized two pairs of Stokeslets for early stage sea star and three pairs of Stokeslets for the late stage sea star larvae under the same experimental confinement conditions () and physical location around the body surface (Methods, Supplementary Information), and showed excellent agreement with the experiments for both types of sea star larvae (Fig. 2a, c, right panels) (Extended Data Fig. 4a, d). It must be noted that in the theoretical model, each Stokeslet generates 1 pair of vortices. However, the resultant velocity field due to the superposition of multiple Stokeslets gave rise to a flow field consisting of two vortices under weak confinement, closely matching the experiments (Extended Data Fig. 4, Methods, Supplementary Information).
@@
In sea urchin larvae, weak confinement () resulted in vortices that are highly elongated in an ellipsoidal shape and distributed over the sides of the arms. Additionally, the vortex centers are located near the tip of the extended arms (Fig. 2e left panel). The theoretical model here utilized two pairs of Stokeslets under the same experimental confinement conditions () and physical location at the tip of the extended arms (Methods, Supplementary Information), and showed a flow field comparable in magnitude with the experiments, but not in the number of vortices generated (Fig. 2e, right panel). Compared to the sea stars, sea urchin morphology is much more complex, with with extended rigid arms. In this developmental stage, the sea urchin larva also had two small rudimentary arms near its oral surface, and these could also potentially influence the observed flow fields. For simplicity, the theoretical model does not account for the presence of these small arms. Hence, the theoretical flow fields are slightly different from that of experiments. Overall, each of the experimental larval vortices under weak confinement can be modeled theoretically with the superposition of multiple Stokeslets, and the experimental two-vortex larval flow field could be captured accurately.
@@
Under strong confinement conditions (), the local morphological variations between the three larvae play an important role in the number of vortices generated (Fig. 2b, d, f, left panels). While the early stage sea star generates two pairs of vortices (Pairs 1, 2) (Fig. 2b), both the late stage sea star and sea urchin larvae generate three pairs of vortices (Pairs 1, 2, 3) (Fig. 2d, f). Hence, strong confinement promotes the development of additional pairs of vortices in all the larvae (Fig. 2b, d, f). The early stage sea star larva has a smooth oval/ellipsoid-shaped body plan, and although it does not yet have arms protruding outside the body, there is a sharp convex curvature change in the region where arms will be formed in the future. The theoretical model with two pairs of Stokeslets placed at exactly the region corresponding to the convex curvature change shows excellent agreement with the experiments (Fig. 2b, right panel).